OpenStudy (anonymous):

The sum of the measures of the interior angles of a regular polygon is 1980. Find the number of sides in the polygon.

OpenStudy (anonymous):

Sum of exterior angles is 360 degrees. Any one exterior angle would be \(\dfrac{360}{n}\) degrees for a regular \(n\)-gon. Exterior angles supplement interior angles, so any one interior angle would be \(\left(180-\dfrac{360}{n}\right)\) degrees. There are \(n\) interior angles to an \(n\)-gon, so the sum would be \((180n-360)\) degrees. \[180n-360=1980\] Solve for \(n\).

OpenStudy (anonymous):

ohhhh thanks bruh